OpenStudy (anonymous):

Add. 5 and 1\8 + 5\6 + 5\12 A. 6 and 3\4 B. 6 and 3\8 C. 6 and 3/12 D. 6 and 3/24

OpenStudy (mathmale):

Hi again, Alex, Learn anything from the previous example that you could apply here? Notice that in the previous example, I wrote\[1 +\frac{ 3 }{ 4 } = \frac{ 4 }{ 4}+\frac{ 3 }{ 4 }=\frac{ 7 }{ 4 }\] Would you mind explaining in your own words what's happening there and why?

OpenStudy (anonymous):

idk i am horrible at math i am only good at science biology and health science

OpenStudy (mathmale):

I don't mean to be sarcastic, but this OpenStudy/Math is your big chance to improve your math skills. Hope you want to. To answer my own question: To combine 1 and 3/4, we need to convert that 1 into a fraction that shares the same denominator with 3/4. Thus, we multiply that 1 by 4/4 and obtain 4/4; then we have 4/4 + 3/4 = 7/4. OK with that? If you are, then please combine 5 and 1/8 (fromthe present problem).

OpenStudy (mathmale):

How's it going, Alex?

OpenStudy (anonymous):

good

OpenStudy (mathmale):

I'm glad. But are you going to try combining 5 and 1/8?

OpenStudy (anonymous):

5 and 1/8 + 5/6 + 5/12 is the same as saying 41/8 + 5/6 + 5/12. Now we need to find a common denominator. The least common multiple of the 8, 6, and 12 is 24. So we after multiplying 41/8 with 3, 5/6 by 4, and 5/12 with 2 to get a common denominator of 24 . We get 123/24 + 20/24 + 10/24. Now we add it all together and get 153/24. Now simplify. 153/24= 6 and 9/24. But we can still simplify it further 6 and 9/24 = 6 and 3/8. You can also leave the 5 just as it is and multiply 1/8 by 3 to get the common denominator of 24. So it would be 5 and 3/24 + 20/ 24 + 10/ 24. When you add the numerators together you get 5 and 33/24. And you simplify it to 6 and 3/8. You can do it either way but leaving the 5 as it is, is easier. I hope you learned!

OpenStudy (mathmale):

Alex: Notice how Princess has correctly obtained 41/8 as the sum of 5 and 1/8. Please, would you try this yourself, so you can be certain of knowing how to do it?

OpenStudy (mathmale):

Alex: Princess goes on to combine 5/6 and 5/12. Do these two fractions have the same or different denominators? If different, what are we to do about that?

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

Its D.

OpenStudy (mathmale):

Mind explaining how you decided that D is the correct answer? I'm always more interested in how you get from A to B than I am in the "answer."

OpenStudy (mathmale):

The Internet is such a wonderful resource. But what if you're taking a test and don't have Internet access during the test? Do you want to learn the technique of doing the problem the traditional way, on paper?